OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6). y = x + 7 x= 1 over 3 y = −3x + 3 y = x + 3 x= 1 over 3 y = −3x + 7

OpenStudy (anonymous):

a line that is parallel would have the same slope, so it would be y=-3x+b

OpenStudy (anonymous):

so all you have to do is take the two point and plug it into that equation and solve for b

OpenStudy (anonymous):

6=-3(-1)+b 6=3+b b=3 the equation would be y=-3x+3